OpenStudy (anonymous):

Consider the function below. f(x)=(x+6)/(x+4) Evaluate the difference quotient for the given function. Simplify your answer. f(x)-f(4)/(x-4)??

jimthompson5910 (jim_thompson5910):

Consider the function below. f(x)=(x+6)/(x+4) Evaluate the difference quotient for the given function. Simplify your answer. f(x)-f(4)/(x-4)?? \[\large f(x)=\frac{x+6}{x+4}\] \[\large f(4)=\frac{4+6}{4+4}\] \[\large f(4)=\frac{10}{8}\] \[\large f(4)=\frac{5}{4}\] ---------------------------------------------------------------------- \[\large \frac{f(x)-f(4)}{x-4}\] \[\large \frac{\frac{x+6}{x+4}-\frac{5}{4}}{x-4}\] \[\large \frac{\frac{4(x+6)}{4(x+4)}-\frac{5(x+4)}{4(x+4)}}{x-4}\] \[\large \frac{\frac{4x+24}{4(x+4)}-\frac{5x+20}{4(x+4)}}{x-4}\] \[\large \frac{\frac{4x+24-(5x+20)}{4(x+4)}}{x-4}\] \[\large \frac{\frac{4x+24-5x-20}{4(x+4)}}{x-4}\] \[\large \frac{\frac{-x+4}{4(x+4)}}{x-4}\] \[\large \frac{-x+4}{4(x+4)}\times\frac{1}{x-4}\] \[\large \frac{-x+4}{4(x+4)(x-4)}\] \[\large \frac{-(x-4)}{4(x+4)(x-4)}\] \[\large \frac{-\cancel{(x-4)}}{4(x+4)\cancel{(x-4)}}\] \[\large \frac{-1}{4(x+4)}\] \[\large -\frac{1}{4x+16}\] So \[\large \frac{f(x)-f(4)}{x-4}=-\frac{1}{4x+16}\]